# Differential Mamba

### v1

@@ A.1 Language Modeling
We train all models with a maximum sequence length of 512 for 40 epochs using 3 seeds (0, 42, 77) with settings as in Table 3. Diff-Mamba uses 1024 channels and 64 heads, while Mamba uses 1024 channels and 128 heads.
Model Dataset # Layers Dropout Batch Size lr
Mamba Wikitext-103 6 0.4 170 5e-4
Diff-Mamba Wikitext-103 6 0.4 170 5e-4
@@ A.2 Ablations Analysis
Table 5: Models architecture details used for the ablation study
@@
A.3 Retrieval
For retrieval experiments, we first convert a Mamba model to Diff-Mamba by finetune it over PG19, as described in Section A.3.1, followed by evaluation on BABILong.
A.3.1 PG19 Fine-tuning
On the retrieval task on BABILong we used Mamba with 370M params pretrained on The Pile for 300B tokens. Since we don’t have sufficient resources to pretrain from scratch Diff-Mamba we aligned the last 12 layers of Mamba model into Diff-Mamba and finetuned on PG19 dataset for 2B tokens. Through the alignment, half the parameters were loaded to $\text{Mamba}_{1}(X)$ and the other half to $\text{Mamba}_{2}(X)$ in each Diff-Mamba layer. This method preserved some of the pretrained weights in those layers. As mentioned, Mamba model was finetuned as well for correct comparison. Both models were finetuned on PG19, lr 3e-4, max length of 2048, and a batch size of  0.5M tokens. Training curves and full configuration are presented in Figure 7 and Table 6.
Figure 7: Training curve of Mamba and Diff-Mamba during the fine-tuning of PG19.
Table 6: Hyperparameters used for the PG19 fine-tuning.
Params Values
Layers 48
lr 3e-4
Max Length 2048
Steps 4000
Batch Size 0.5M tokens
Warmup Steps 400
Optimizer AdamW
Weight Decay 0.1
The high loss at the beginning of training Diff-Mamba is due to the alignment process, which is not perfect, and the adaptation process of the new architecture. The training is not sufficient and more steps should be done for full adaptation. Nonetheless, Diff-Mamba demonstrates impressive capabilities even in the current circumstances. For more clearer presentation, we show original scores for each model in Figure 8.
(a) Zero-Shot Diff-Mamba
(b) Zero-Shot Mamba
Figure 8: Comparison of original scores of Diff-Mamba and Mamba models as Zero-Shot. Values are the percentage of correct examples while green indicates larger score than the other model and red the opposite.
A.3.2 BABILong Fine-tuning
The models were fine-tuned on 90% examples up to 1k tokens inside the dataset across all tasks, which is approximately equivalent to 17k examples. The other 10% remained to test. Following [22], we employ a preprocessing step that shapes each sample as follows: "<context>{input}</context>Question:{question} Answer:" and the loss was calculated on the answer label only. Since the training tiny size, three seeds (0, 42, 77) have been tested and averaged through the results. Original scores for each model are presented in Figure 9. The configuration for the training is in Table 7.
Table 7: Hyperparameters used for the BABILong fine-tuning
@@ A.3.2 BABILong Fine-tuning
(a) Fine-tuned Diff-Mamba
(b) Fine-tuned Mamba
Figure 9: Comparison of original scores of Diff-Mamba and Mamba models after fine-tuning on BABILong. Values are the percentage of correct examples while green indicates larger score than the other model and red the opposite.
A.4 Noise Analysis in Intermediate Representations
To measure the signal-to-noise ratio, we employ Tuned-Lens [4]. Specifically, we train linear projections to learn the appropriate transformation over intermediate representations across layers. The lens for finetuned BABILong models have been trained on the validation set of The Pile with a similar configuration to the original tuned lens paper. Evaluation has been done on the test set created in the retrieval task, which stands for 2k examples. For each example, the needle probability is gathered from each layer.
A.5 Ablations for Early Results at Medium Scale
We pre-trained both the Mamba and Diff-Mamba architectures and conducted an ablation study on the Diff-Mamba model to evaluate the impact of architectural variations. The results indicate that interleaving Mamba and Diff-Mamba layers throughout the network yields improved loss performance during pre-training, compared to using a fully Diff-Mamba architecture (Figure 10). The pre-training was performed on 8 GPUs over 50 billion tokens using the GPTNeoX tokenizer. The Mamba model contains 368 million parameters, while the Diff-Mamba-Alternate and Diff-Mamba-Full configurations comprise 375 million and 382 million parameters, respectively. All models were trained with 48 layers, a learning rate of 1.5e-3, 10,000 warm-up steps, the AdamW optimizer with a weight decay of 0.1, a batch size of 1 million tokens per step, and a maximum sequence length of 2048 tokens.
Figure 10: Loss training curve of Diff-Mamba variants compared to Mamba through pre-training on The-Pile. Diff-Mamba-Alternate indicates alternating layers of Mamba and Diff-Mamba, while Diff-Mamba-Full indicates only Diff-Mamba layers model. The trends were smoothed for display adjustment.

### v2

@@ A.1 Language Modeling
We train all models with a maximum sequence length of 512 for 40 epochs using 3 seeds (0, 42, 77) with settings as in Table 3. Diff-Mamba uses 1024 channels and 64 heads, while Mamba uses 1024 channels and 128 heads. In Diff-Mamba, we reduce the number of parameters in the linear projections of each Mamba block by half to achieve a similar number of parameters to Mamba. Following (Ye et al., 2024), the parameter $\lambda_{\text{init}}$ in Eq. 11 is defined as $\lambda_{\text{init}}=0.8-0.6\cdot\exp\!\bigl(-0.3\,(i_{\text{layer}}-1)\bigr)$. In addition, the normalization we use in Eq. 21 is RMS-Norm (Zhang and Sennrich, 2019).
Model Dataset # Layers Dropout Batch lr
Mamba Wikitext-103 6 0.25 100 5e-4
Diff-Mamba Wikitext-103 6 0.25 100 5e-4
@@ A.2 Ablations Analysis
Table 4: Hyperparameters used for the ablation trainings.
@@
Table 5: Models architecture details used for the ablation study.
Figure 7: Loss training curve of Diff-Mamba variants compared to Mamba through pre-training on The Pile. Diff-Mamba-Hybrid indicates alternating layers of Mamba and Diff-Mamba, while Diff-Mamba-Full indicates only Diff-Mamba layers model. The trends were smoothed for display adjustment.
A.3 Unit Tests as Predictors of Scaling
We follow the MAD pipeline Poli et al. (2024) and adopt a rigorous suite of synthetic token manipulation tasks, that serve as capability unit tests. We investigate Mamba and Diff-Mamba. Each model has four layers, consisting of the core blocks. Hyperparameters and a set of tests were applied similarly to the MAD pipeline, including different numbers of examples for training, learning rates, sequence lengths, and vocabulary size upon tasks such as compression, ICR, and selective-copying.
A.4 Early Results at Medium Scale
We pre-trained both the Mamba and Diff-Mamba architectures and conducted an ablation study on the Diff-Mamba model to evaluate the impact of architectural variations. The results indicate that interleaving Mamba and Diff-Mamba layers throughout the network yeilds improved loss performance during pre-training, compared to using fully Diff-Mamba architecture (Figure 7). The pre-training was performed on 8 GPUs over 50 billion tokens using the GPTNeoX tokenizer. The Mamba model contains 368M parameters, Diff-Mamba-Hybrid has 375M parameters, and Diff-Mamba-Full configuration comprises 382M parameters. Although Diff-Mamba-Hybrid has 7 million fewer parameters than Diff-Mamba-Full, it achieves better performance. The models were trained with 48 layers, a learning rate of 1.5e-3, 10,000 warm-up steps, the AdamW optimizer with a weight decay of 0.1, a batch size of 1M tokens per step, and a maximum sequence length of 2048 tokens.
A.5 Retrieval
For retrieval experiments, we use 50B checkpoint of Diff-Mamba and Mamba trained on The Pile, as described in Section 4.4 and finetune those on BABILong tasks, followed by an evaluation on a test set. For statistically significant results, we finetuned and evaluated with 3 seeds.
A.5.1 BABILong Fine-tuning
The models were fine-tuned on 90% examples up to 1k tokens inside the dataset across all tasks, which is approximately equivalent to 17k examples. The other 10% remained for test. Following (Kuratov et al., 2024), we employ a preprocessing step that shapes each sample as follows: "<context>{input}</context> Question:{question} Answer:" and the loss was calculated on the answer label only. Since the training tiny size, three seeds have been tested and averaged through the results. Original scores for each model are presented in Figure 8. The configuration for the training is in Table 6.
@@ A.5.1 BABILong Fine-tuning
Table 6: Hyperparameters used for the BABILong fine-tuning
(a) Diff-Mamba
(b) Mamba
Figure 8: Comparison of original scores of Diff-Mamba and Mamba models after fine-tuning on BABILong. Values are the percentage of correct examples, while green indicates a larger score than the other model, and red indicates the opposite.
A.6 Noise in Intermediate Representations
To measure the signal-to-noise ratio, we employ Tuned-lens (Belrose et al., 2023). Specifically, we train linear projections to learn the appropriate transformation over intermediate representations across layers and we demonstrate it using the BABILong finetuned models from Section 4.5. The probe for finetuned BABILong models have been trained on the validation set of The Pile for 3 seeds with a similar configuration to the original Tuned-lens paper. Evaluation was conducted on a test set derived from the BABILong benchmark, consisting of 1,000 examples with context lengths ranging from 1k to 2k tokens. These examples were sampled from the first five tasks, with 200 examples per task. For each example, the needle probability was extracted from every layer.
Appendix B Efficiency Benchmark
We evaluate the computational efficiency of Diff-Mamba-Hybrid (375M params) compared to Mamba (368M params) with respect to inference speed, memory footprint, and forward-pass latency. These measurements provide a comprehensive view of the trade-offs introduced by the architectural modifications.
B.1 End-to-End Inference
We calculate end-to-end inference throughput across varying batch sizes, measured on an L40S GPU with 48GB of memory (Figure 9). The benchmark was conducted using a prompt length of 2048 tokens and generating 128 new tokens with untrained models of Mamba, Diff-Mamba-Full, Diff-Mamba-Hybrid, and Transformer for a baseline. Throughput (tokens/s) is computed as batch size $\times$ 128 / inference time. Both fully and hybrid Diff-Mamba exhibits lower throughput compared to the original Mamba at small batch sizes; however, this gap narrows considerably as the batch size increases, and becomes only a 12% difference with a batch size of 64 for Diff-Mamba-Hybrid.
B.2 Memory Benchmark
GPU memory usage across different batch sizes is reported in Table 7. Each batch consists of sequences of length 2048, and memory was measured after a forward pass. Diff-Mamba incurs only a modest increase in memory consumption relative to Mamba, with a difference of less than 1GB even at batch size 32 (44.30GB vs. 43.80GB). This result suggests that the additional computational layers introduced in Diff-Mamba achieve efficiency improvements without imposing prohibitive memory demands, thereby maintaining practical feasibility in resource-constrained environments.
Batch Size Mamba (GB) Diff-Mamba (GB)
1 3.92 3.95
2 5.16 5.25
4 7.72 7.84
8 12.84 13.03
16 23.09 23.43
32 43.80 44.30
Table 7: Memory usage comparison across different batch sizes.
B.3 Forward-Pass Latency
We further examine forward-pass latency across varying sequence lengths, summarized in Table 8. Forward-pass time serves as a direct proxy for training efficiency, since it closely aligns with per-step training duration. The results demonstrate that Diff-Mamba sustains comparable per-step efficiency to Mamba, introducing only marginal latency overhead at extended sequence lengths.
Figure 9: Inference throughput on L40s 48GB (prompt length 2048).
Seq Length Mamba (s) Diff-Mamba (s)
512 0.063 0.076
1024 0.079 0.090
2048 0.073 0.087
4096 0.147 0.172
8192 0.271 0.282
Table 8: Forward time comparison for different sequence lengths.